Mr. Thomas invested an amount of Rs. 13,900 divided in two different schemes A and B at the simple interest rate of 14% p.a. and 11% p.a. respectively. If the total amount of simple interest earned in 2 years be Rs. 3508, what was the amount invested in Scheme B?
1
Expert's answer
2019-04-09T09:20:57-0400

Sum scheme A -  x

Sum scheme B - (13900 - x).

Then,

((14*2*x)/100)+((13900-x)*11*2)/100=3508

28x-22x=350800-(13900*22)

6x=45000

x=7500 – in scheme A

13900-7500=6400 – in scheme B.